OpenStudy (anonymous):

The table below represents the distance a person runs as a function of time: The average rate of change of the function between x = 10 to x = 16 is __________ m/s and represents the person's speed.

Nnesha (nnesha):

where is dining table

OpenStudy (anonymous):

sorry os crashed http://goo.gl/pR4oBB

OpenStudy (anonymous):

i say 3, wouldn't it?

OpenStudy (anonymous):

wait, average rate of change would be the slope, right?

OpenStudy (anonymous):

average rate of change = f(b) - f(a)/b - a right?

Nnesha (nnesha):

yes right

Nnesha (nnesha):

nope 3 is not right :(

OpenStudy (agl202):

average rate of change = (change in y)/(change in x) = (96 - 60) / (16 - 10) = 36/6 = 6 Hope this helped!

OpenStudy (anonymous):

ok so avergae rate = 96 - 60 / 16 - 10 = 36/6 = 6

Nnesha (nnesha):

lO_Oks gO_Od :)

OpenStudy (agl202):

yep

Nnesha (nnesha):

wait what hmm i didn't see that u already got the answer :/ nvm :/

OpenStudy (anonymous):

oh I didnt even see her reply until you said that